OpenStudy (anonymous):

Can anyone help me? 8-4y=-7-7y

OpenStudy (angela210793):

8-4y+7+7y=0 3y+a5=0 3y=-15 y=-15/3=-5 were u looking this?

OpenStudy (anonymous):

Yes thank you